# Differentially Private Graph Diffusion with Applications in Personalized PageRanks

### v1

@@ Abstract
Graph diffusion, which iteratively propagates real-valued substances among the graph, is used in numerous graph/network-involved applications. However, releasing diffusion vectors may reveal sensitive linking information in the data such as transaction information in financial network data. However, protecting the privacy of graph data is challenging due to its interconnected nature. This work proposes a novel graph diffusion framework with edge-level different privacy guarantees by using noisy diffusion iterates. The algorithm injects Laplace noise per diffusion iteration and adopts a degree-based thresholding function to mitigate the high sensitivity induced by low-degree nodes. Our privacy loss analysis is based on Privacy Amplification by Iteration (PABI), which to our best knowledge, is the first effort that analyzes PABI with Laplace noise and provides relevant applications. We also introduce a novel $\infty$-Wasserstein distance tracking method, which tightens the analysis of privacy leakage and makes PABI more applicable in practice. We evaluate this framework by applying it to Personalized Pagerank computation for ranking tasks. Experiments on real-world network data demonstrate the superiority of our method under stringent privacy conditions.
@@ 1.1 More Related Works
In the realm of graph diffusion, the most related work to ours is by Epasto et al. [20], which focuses on releasing the PPR vector using forward push [41] and Laplace output perturbation. Some studies have shown that injecting noise during the process may offer better privacy-utility trade-offs compared to output perturbation methods [25] and our study compared with [20] provides another use case. Traditionally, the composition theorem [42] is used to track privacy guarantees for iterative algorithms, but it results in a bound that may diverge with the number of iterations. Recently, the technique of PABI [26, 27] was introduced to strengthen the privacy analysis of adding noise during the process, which demonstrates a non-divergent privacy bound for releasing final results if the iterations adopt contraction maps [27]. This substantially tightens the divergent bound given by the naive application of the DP composition theorem [43, 44]. Our framework also benefits from this advantage and we further adapt Altschuler et al.’s analysis [27] to incorporate the Laplace mechanism and provide a tightened space diameter tracking, which makes the bound practically applicable in the graph diffusion application.
@@ 2 Preliminaries
where $\mathbf{s}\in\mathbb{R}^{|\mathcal{V}|}$ is a stochastic vector on the graph, and $\gamma_{1,k}+\gamma_{2,k}+\gamma_{3,k}=1$. Let $\gamma_{\text{max}}=\max_{k}|\gamma_{1,k}|+|\gamma_{2,k}|$ denote the Lipschitz constant of the graph diffusion mapping. $\mathbf{s}_{K}$ is the diffusion vector at time $K$. The essence of a diffusion process is to model how an initial vector $\mathbf{s}$ propagates through the graph over time. Coefficient $\gamma_{i,k}$’s control how different resources contribute to the $k$th step diffusion. When taking $\gamma_{1,k}=1-\gamma_{3,k}=\beta$ and $\gamma_{2,k}=0$, Eq. (1) is recognized as the PageRank Diffusion [46] with teleport probability $1-\beta$. The Exponential kernel diffusion, which includes the specific case of the Heat Kernel diffusion [2], can also be characterized with the composition of diffusion mappings via the infinitely divisible property [47].
@@ 3 Methodology
This study centers on a category of $\gamma_{\text{max}}<1$ Lipschitz continuous graph diffusions, encompassing prevalent techniques such as PageRank [1] and PPR [53]. It is noted that each diffusion map $\phi_{k}$ within graph diffusion maintains the total probability mass of the diffusion vector $\mathbf{s}_{k}$, owing to the condition $\sum_{i\in[3]}\gamma_{i,k}=1$ and the property of the random walk matrix $\mathbf{P}$ being a left stochastic matrix. This observation entails that the diffusion map $\phi_{k}$ in Eq. (1) constitutes a strictly contraction map in the metric space $(\mathbb{R}^{|\mathcal{V}|},\|\cdot\|_{1})$. Consequently, graph diffusion can be construed as a composite of contraction maps. The PABI technique has been devised to privatize contractive iterations by injecting random noise per iteration. Empirical studies suggest that distributing noise throughout the diffusion steps can provide improved utility-privacy trade-offs compared to output perturbation alone [25]. This insight serves as a key motivation for employing PABI to establish privacy-preserving graph diffusion.
@@ 3.2 Private Graph Diffusions
where $f$ is a graph-dependent degree-based function set as $f(\mathbf{x})=\min(\max(\mathbf{x},0),\eta\cdot\mathbf{d})$ for $\phi_{k,\sigma}$, Noise variables $\xi_{k}^{(1)}$ and $\xi_{k}^{(2)}$ are independently sampled from a Laplace distribution $\mathcal{L}(\mathbf{0},\sigma)$. It is noteworthy that our framework can also be extended to accommodate Gaussian distributions. However, Gaussian noise has been shown to be suboptimal for graph diffusion in $\ell_{1}$ space, with empirical evidence provided in App. D.4.
@@
$\epsilon\leq\min_{\tau\in\{0,1,...,K-1\}}\left[(K-\tau)\cdot g_{\alpha}\left(\sigma,\rho_{\text{diff}}\right)+g_{\alpha}\left(\sigma,\frac{\rho_{\text{diff}}}{1-\gamma_{\text{max}}}\cdot\gamma_{\text{max}}^{K-\tau}\right)\right]$ (5)
where $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho))$ denotes the Rényi divergence induced by the Laplace mechanism [44]. The distortion $\rho_{\text{diff}}=(4\gamma_{\text{max}}^{(1)}+2\gamma_{\text{max}})\eta$ involves Lipschitz continuity coefficient $\gamma_{\text{max}}$, and maximum diffusion coefficient $\gamma_{\text{max}}^{(1)}=\max_{k\in[K]}|\gamma_{1,k}|$.
@@
The privacy bound in Eq. (5) consists of two components: the distortion absorption term (the first term on the RHS) and the PABI term (the second term in RHS). Distortion absorption quantifies the cumulative Rényi divergence over Laplace noise with distortion $\rho_{\text{diff}}$, while the PABI term quantifies the exponential decay rate, echoing the result in Eq. (3). However, a key difference lies in our approach; instead of leveraging the projected set diameter $D$ to control the distance between coupled CNIs, our proposed $\infty$-Wasserstein tracking method yields a more practical term, $\frac{\rho_{\text{diff}}}{1-\gamma_{\text{max}}}$. Further details and utility evaluations of this tool are presented in the proof sketch and Sec. 4.2, respectively.
@@
In Eq. (6), we demonstrate the convergence of the privacy budget with respect to diffusion steps $K$. This approach differs from the adaptive composition theorem [42], which analyzes how privacy guarantees degrade when composed mechanisms are applied. Although this method has commonly been employed to protect privacy in graph learning models [50, 54, 51], it leads to a linear increase in the privacy budget with the number of iterations $K$ under Rényi divergence [44], potentially resulting in unbounded losses as $K$ grows to infinity.
Figure 2: RDP vs. Total Diffusion Step $K$ with $\eta=10^{-5}$.
More importantly, even for a small number of diffusion steps, our framework achieves a significantly better privacy budget under practical PPR diffusion settings, as illustrated in Fig. 2. Further empirical evaluations are detailed in App. D.4.
Note that our framework employs a uniform noise scale throughout the diffusion process. However, our analysis can adapt to varied noise allocations, as detailed in App. B. This flexibility supports potentially better utility under specific privacy constraints.
@@ 3.3 Proof Sketch of Theorem 1
Here , $\tilde{\rho}=4\gamma_{\text{max}}^{(1)}\cdot\max_{\mathbf{x}\in\mathbb{R}^{n}}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|+\gamma_{\text{max}}\cdot\Delta f$ measures the distortion incurred by single step diffusion, where $\Delta f=\sup_{\mathbf{x}\in\mathbb{R}^{n}}\|f(\mathbf{x})-f^{\prime}(\mathbf{x})\|_{1}$.
The distortion $\tilde{\rho}$ depends on the thresholding function $f$ normalized by node degrees, i.e., $\max_{\mathbf{x}\in\mathbb{R}^{n}}\left|\frac{[f(\mathbf{x})]_{i}}{d_{i}}\right|$. This observation highlights the importance of a degree-based design for the thresholding function $f$. Uniform thresholding across all nodes results in distortion proportional to $\frac{1}{d_{\text{min}}}$, introducing unnecessarily large noise induced by low-degree nodes and degrading overall performance. This in principle inspires the choice of $f$ relying on node degrees. Consequently, $\tilde{\rho}$ is bounded by $\rho_{\text{diff}}=(4\gamma_{1,\text{max}}+2\gamma_{\text{max}})\eta$.
Figure 3: Setting: Graph Diffusion with $\gamma_{1,k}=0.8,\gamma_{3,k}=0.2$.
Step 4: Upper Bounding PABI with $\bm{\infty}$-Wasserstein distance tracking. To perform tight privacy analysis for the second term in Eq. (7), we develop a novel $\infty$-Wasserstein distance tracking method for coupled CNIs. This method discards the original boundedness condition in PABI (Eq. (3) Second Term), which relies on the diameter $D$.
@@
Given two coupled graph diffusions mentioned above, for any $\tau\geq 0$, any noise realization $\zeta$, we have
$\mathcal{D}_{\alpha}(\mathbf{s}_{K}|\xi_{\tau+1:K}^{(2)}=\zeta\|\mathbf{s}_{K}^{\prime}|\tilde{\xi}_{\tau+1:K}^{\prime(2)}=\zeta)\leq g_{\alpha}(\sigma,\gamma_{\text{max}}^{K-\tau}w)$ (9)
where $w=\frac{\rho_{\text{diff}}}{1-\gamma_{\text{max}}}$ represents the tracked $\infty$-Wasserstein distance over coupled CNIs.
We argue that using $w$ instead of the default diameter $D$ is crucial to make the algorithm practically useful. There is no numerical evaluation in the previous study [27]. Numerical comparison between $w$ and the diameter of thresholding function $D=\eta\cdot\sum_{i=1}^{|\mathcal{V}|}d_{i}$, using the real-world BlogCatalog dataset (detailed in Sec. 4), is illustrated in Fig.3. $w$ achieves orders-of-magnitude improvement, which is still significant even if $D$ impacts privacy loss via a logarithmic term. Further empirical validations demonstrating significant utility improvements are detailed in Sec. 4.2.
@@ 3.4 Personalized Graph Diffusion Algorithms with Application in PPR Diffusion
In practice, graph diffusions often originate from a single node $\mathbf{e}_{i}$, personalizing the algorithm to this seed node (user). Since the output is provided only to the seed node, protecting its edge connections becomes unnecessary, ensuring no privacy leakage in the first diffusion step under personalized privacy guarantees. Consequently, the thresholding function is tailored as follows: $f(\mathbf{x})=\min(\max(\mathbf{x},0),\eta\cdot\tilde{\mathbf{d}})$, where $[\tilde{\mathbf{d}}]_{j}=[\mathbf{d}]_{j}$ for $j\neq i$ and $[\tilde{\mathbf{d}}]_{i}$ can be set to any positive threshold. We employ personalized edge-level RDP (Definition 2), caring two adjacent graphs with a difference only in a single edge not linked directly to the seed node. This approach is encapsulated in the following theorem:
@@
Given a graph $\mathcal{G}$, an associate graph diffusion $\mathscr{D}=\{\phi_{k}\}_{k=1}^{\infty}$, then personalized noisy graph diffusion mechanism $\mathscr{D}_{K,\sigma}$ with corresponding $f(\mathbf{x})=\min(\max(\mathbf{x},0),\eta\cdot\tilde{\mathbf{d}})$ ensures personalized edge-level $(\alpha,\epsilon)$-RDP with $\epsilon$ satisfies:
$\epsilon\leq\min_{\tau\in\{0,1,...,K-1\}}\left[g_{\alpha}\left(\sigma,\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})\cdot\gamma_{\text{max}}^{K-\tau}}{1-\gamma_{\text{max}}}\right)+(K-\tau)\cdot g_{\alpha}\left(\sigma,\rho_{\text{diff}}\cdot\mathbbm{1}_{\tau\neq 0}\right)\right]$ (10)
@@
where lazy random walk matrix $\mathbf{W}=\frac{1}{2}(\mathbf{P}_{\text{rw}}+\mathbf{I})$ and $(1-\beta)$ represents teleport probability.
Our framework incorporates noise into the diffusion process of each step of PPR. The privacy guarantees for this noisy PPR are derived from Theorem 4 with $\rho_{\text{diff}}=4\beta\eta$ and $\gamma_{\text{max}}=\beta$.

### v4

@@ Abstract
Graph diffusion, which iteratively propagates real-valued substances among the graph, is used in numerous graph/network-involved applications. However, releasing diffusion vectors may reveal sensitive linking information in the data such as transaction information in financial network data. However, protecting the privacy of graph data is challenging due to its interconnected nature. This work proposes a novel graph diffusion framework with edge-level differential privacy guarantees by using noisy diffusion iterates. The algorithm injects Laplace noise per diffusion iteration and adopts a degree-based thresholding function to mitigate the high sensitivity induced by low-degree nodes. Our privacy loss analysis is based on Privacy Amplification by Iteration (PABI), which to our best knowledge, is the first effort that analyzes PABI with Laplace noise and provides relevant applications. We also introduce a novel $\infty$-Wasserstein distance tracking method, which tightens the analysis of privacy leakage and makes PABI more applicable in practice. We evaluate this framework by applying it to Personalized Pagerank computation for ranking tasks. Experiments on real-world network data demonstrate the superiority of our method under stringent privacy conditions.
@@ 1.1 More Related Works
In the realm of graph diffusion, the most related work to ours is by Epasto et al. [20], which focuses on releasing the PPR vector using forward push [41] and Laplace output perturbation. Some studies have shown that injecting noise during the process may offer better privacy-utility trade-offs compared to output perturbation methods [25] and our study compared with [20] provides another use case. Traditionally, the composition theorem [42] is used to track privacy guarantees for iterative algorithms, but it results in a bound that may diverge with the number of iterations. Recently, the technique of PABI [26, 27] was introduced to strengthen the privacy analysis of adding noise during the process, which demonstrates a non-divergent privacy bound for releasing final results if the iterations adopt contraction maps [27]. This substantially tightens the divergent bound given by the naive application of the DP composition theorem [43, 44]. Our framework also benefits from this advantage and we further adapt Altschuler et al.’s analysis [27] to incorporate the Laplace mechanism and provide a tightened space diameter tracking, which makes the bound practically applicable in the graph diffusion application. Besides, works that share a similar spirit in leveraging PABI have been conducted for other scenarios including machine unlearning [45] and improving hidden state DP [46].
@@ 2 Preliminaries
where $\mathbf{s}\in\mathbb{R}^{|\mathcal{V}|}$ is a stochastic vector on the graph, and $\gamma_{1,k}+\gamma_{2,k}+\gamma_{3,k}=1$. Let $\gamma_{\text{max}}=\max_{k}|\gamma_{1,k}|+|\gamma_{2,k}|$ denote the Lipschitz constant of the graph diffusion mapping, and $\gamma_{\text{max}}^{(1)}=\max_{k}|\gamma_{1,k}|$ denote the maximum diffusion coefficient. $\mathbf{s}_{K}$ is the diffusion vector at time $K$. The essence of a diffusion process is to model how an initial vector $\mathbf{s}$ propagates through the graph over time. Coefficient $\gamma_{i,k}$’s control how different resources contribute to the $k$th step diffusion. When taking $\gamma_{1,k}=1-\gamma_{3,k}=\beta$ and $\gamma_{2,k}=0$, Eq. (1) is recognized as the PageRank Diffusion [48] with teleport probability $1-\beta$. The Exponential kernel diffusion, which includes the specific case of the Heat Kernel diffusion [2], can also be characterized with the composition of diffusion mappings via the infinitely divisible property [49].
@@ 3 Methodology
This study centers on a category of $\gamma_{\text{max}}<1$ Lipschitz continuous graph diffusions, encompassing prevalent techniques such as PageRank [1] and PPR [55]. It is noted that each diffusion map $\phi_{k}$ within graph diffusion maintains the total mass of the diffusion vector $\mathbf{s}_{k}$, owing to the condition $\sum_{i\in[3]}\gamma_{i,k}=1$ and the property of the random walk matrix $\mathbf{P}$ being a left stochastic matrix. This observation entails that the diffusion map $\phi_{k}$ in Eq. (1) constitutes a strictly contraction map in the metric space $(\mathbb{R}^{|\mathcal{V}|},\|\cdot\|_{1})$. Consequently, graph diffusion can be construed as a composite of contraction maps. The PABI technique has been devised to privatize contractive iterations by injecting random noise per iteration. Empirical studies suggest that distributing noise throughout the diffusion steps can provide improved utility-privacy trade-offs compared to output perturbation alone [25]. This insight serves as a key motivation for employing PABI to establish privacy-preserving graph diffusion.
@@ 3.2 Private Graph Diffusions
where $f$ is a graph-dependent degree-based function set as $f(\mathbf{x})=\min(\max(\mathbf{x},-\eta\cdot\mathbf{d}),\eta\cdot\mathbf{d})$ with a threshold parameter $\eta$ to balance privacy-utility trade-off. Specifically, $f$ clips the values of the diffusion vector according to node degrees. Notably, the thresholding function $f$ allows for negative signals, capturing scenarios where the diffusion coefficient $\gamma_{1,k}$ can be negative. Noise variables $\xi_{k}^{(1)}$ and $\xi_{k}^{(2)}$ are independently sampled from a Laplace distribution $\mathcal{L}(\mathbf{0},\sigma)$. It is noteworthy that our framework can also be extended to accommodate Gaussian distributions. However, Gaussian noise has been shown to be suboptimal for graph diffusion in $\ell_{1}$ space, with empirical evidence provided in App. D.4.
@@
$\epsilon\leq\min_{\tau\in\{0,1,...,K-1\}}\left[(K-\tau)\cdot g_{\alpha}\left(\sigma,\rho_{\text{diff}}\right)+g_{\alpha}\left(\sigma,\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})}{1-\gamma_{\text{max}}}\cdot\gamma_{\text{max}}^{K-\tau}\right)\right]$ (5)
where $g_{\alpha}(\sigma,\rho)=\frac{1}{\alpha-1}\ln(\frac{\alpha}{2\alpha-1}\exp(\frac{\alpha-1}{\sigma}\rho)+\frac{\alpha-1}{2\alpha-1}\exp(-\frac{\alpha}{\sigma}\rho))$ denotes the Rényi divergence induced by the Laplace mechanism [44], and $\rho_{\text{diff}}=\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$ represents the maximum single-step distortion incurred by diffusion on adjacent graphs that involves Lipschitz continuity coefficient $\gamma_{\text{max}}$, and maximum diffusion coefficient $\gamma_{\text{max}}^{(1)}$.
@@
The privacy bound in Eq. (5) consists of two components: the distortion absorption term (the first term on the RHS) and the PABI term (the second term in RHS). Distortion absorption quantifies the cumulative Rényi divergence over Laplace noise with single-step distortion $\rho_{\text{diff}}$, while the PABI term quantifies the exponential decay rate, echoing the result in Eq. (3). However, a key difference lies in our approach; instead of leveraging the projected set diameter $D$ to control the distance between coupled CNIs, our proposed $\infty$-Wasserstein tracking method yields a more practical term, $\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})}{1-\gamma_{\text{max}}}$. Further details and utility evaluations of this tool are presented in the proof sketch and Sec. 4.2, respectively.
@@
It is important to note that the maximum single-step distortion $\rho_{\text{diff}}$ in Eq. (5) is tight and conveys several messages. First, as defined in Eq. (1), when the diffusion process is relatively slow (i.e., $\gamma_{1,k}<\gamma_{2,k}$), the distortion remains tight, governed by the Lipschitz constant $\gamma_{\text{max}}$ of the diffusion mapping. In contrast, when the diffusion is relatively fast (i.e., $\gamma_{1,k}\geq\gamma_{2,k}$), the distortion bound becomes asymptotically tight, depending on graph structures, with worst-case scenarios detailed in App. B.1.
Figure 2: RDP vs. Total Diffusion Step $K$ with $\gamma_{1,k}=0.8,\gamma_{2,k}=0,\gamma_{3,k}=0.2,\alpha=2,\sigma=0.01$, and $\eta=10^{-5}$.
In Eq. (6), we demonstrate the convergence of the privacy budget with respect to diffusion steps $K$. This approach differs from the adaptive composition theorem [42], which analyzes how privacy guarantees degrade when composed mechanisms are applied. Although this method has commonly been employed to protect privacy in graph learning models [52, 56, 53], it leads to a linear increase in the privacy budget with the number of iterations $K$ under Rényi divergence [44], potentially resulting in unbounded losses as $K$ grows to infinity. More importantly, even for a small number of diffusion steps, our framework achieves a significantly better privacy budget under practical PPR diffusion settings, as illustrated in Fig. 2. Further empirical evaluations are detailed in App. D.4.
@@ 3.3 Proof Sketch of Theorem 1
Here, $\tilde{\rho}$ quantifies the maximum distortion introduced by a single-step diffusion and is determined by the thresholding function $f$ normalized by node degrees, i.e., $\frac{[f(\mathbf{s}_{k})]_{i}}{d_{i}}$.
The observation on $\tilde{\rho}$ highlights the importance of a degree-based design for the thresholding function $f$. Uniform thresholding across all nodes results in distortion proportional to $\frac{1}{d_{\text{min}}}$, introducing unnecessarily large noise induced by low-degree nodes and degrading overall performance. This in principle inspires the choice of $f$ relying on node degrees. Consequently, $\tilde{\rho}$ is tightly bounded by $\rho_{\text{diff}}=\max(4\gamma_{\text{max}}^{(1)},2\gamma_{\text{max}})\cdot\eta$.
Figure 3: Setting: Graph Diffusion with $\gamma_{1,k}=0.8,\gamma_{2,k}=0,\gamma_{3,k}=0.2$.
Step 4: Upper Bounding PABI with $\bm{\infty}$-Wasserstein distance tracking. To perform tight privacy analysis for the second term in Eq. (7), we develop a novel $\infty$-Wasserstein distance tracking method for coupled CNIs, where we denote the $\infty$-Wasserstein distance at step $\tau$ by $w_{\tau}$. This method discards the original boundedness condition in PABI (Eq. (3) Second Term), which relies on the diameter $D$.
@@
Given two coupled graph diffusions mentioned above, for any $\tau\in\{0,1,...,K-1\}$, any noise realization $\zeta$, we have
$\mathcal{D}_{\alpha}(\mathbf{s}_{K}|\xi_{\tau+1:K}^{(2)}=\zeta\|\mathbf{s}_{K}^{\prime}|\tilde{\xi}_{\tau+1:K}^{\prime(2)}=\zeta)\leq g_{\alpha}(\sigma,\gamma_{\text{max}}^{K-\tau}w_{\tau})$ (9)
where the tracked $\infty$-Wasserstein distance over coupled CNIs is given by $w_{\tau}=\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})}{1-\gamma_{\text{max}}}$ and is naturally upper bounded by $\frac{\rho_{\text{diff}}}{1-\gamma_{\text{max}}}:=w$.
We argue that using $w_{\tau}$ (or the upper bound $w$) instead of the default diameter $D$ is crucial to make the algorithm practically useful. There is no numerical evaluation in the previous study [27]. Numerical comparison between $w$ and the diameter of thresholding function $D=\eta\cdot\sum_{i=1}^{|\mathcal{V}|}d_{i}$, using the real-world BlogCatalog dataset (detailed in Sec. 4), is illustrated in Fig.3. $w$ achieves orders-of-magnitude improvement, which is still significant even if $D$ impacts privacy loss via a logarithmic term. Further empirical validations demonstrating significant utility improvements are detailed in Sec. 4.2.
@@ 3.4 Personalized Graph Diffusion Algorithms with Application in PPR Diffusion
In practice, graph diffusions often originate from a single node $\mathbf{e}_{i}$, personalizing the algorithm to this seed node (user). Since the output is provided only to the seed node, protecting its edge connections (one-hop neighbors) becomes unnecessary, ensuring no privacy leakage in the first diffusion step under personalized privacy guarantees. Consequently, the thresholding function is tailored as follows: $f(\mathbf{x})=\min(\max(\mathbf{x},-\eta\cdot\tilde{\mathbf{d}}),\eta\cdot\tilde{\mathbf{d}})$, where $[\tilde{\mathbf{d}}]_{j}=[\mathbf{d}]_{j}$ for $j\neq i$ and $[\tilde{\mathbf{d}}]_{i}$ can be set to any positive threshold, i.e., no control is needed for the diffusion over seed node. We employ personalized edge-level RDP (Definition 2), caring two adjacent graphs with a difference only in a single edge not linked directly to the seed node. This approach is encapsulated in the following theorem:
@@
Given a graph $\mathcal{G}$, an associate graph diffusion $\mathscr{D}=\{\phi_{k}\}_{k=1}^{\infty}$, then personalized noisy graph diffusion mechanism $\mathscr{D}_{K,\sigma}$ with corresponding $f(\mathbf{x})=\min(\max(\mathbf{x},-\eta\cdot\tilde{\mathbf{d}}),\eta\cdot\tilde{\mathbf{d}})$ ensures personalized edge-level $(\alpha,\epsilon)$-RDP with $\epsilon$ satisfies:
$\epsilon\leq\min_{\tau\in\{0,1,...,K-1\}}\left[(K-\tau)\cdot g_{\alpha}\left(\sigma,\rho_{\text{diff}}\cdot\mathbbm{1}_{\tau\neq 0}\right)+g_{\alpha}\left(\sigma,\frac{\rho_{\text{diff}}\cdot(1-\gamma_{\text{max}}^{\tau})}{1-\gamma_{\text{max}}}\cdot\gamma_{\text{max}}^{K-\tau}\right)\right]$ (10)
@@
Note that a key difference from Theorem 1 is that in personalized privacy settings, there is no privacy leakage in the first diffusion step ($K=1$).
@@
where lazy random walk matrix $\mathbf{W}=\frac{1}{2}(\mathbf{P}+\mathbf{I})$ and $(1-\beta)$ represents teleport probability with $\beta\in(0,1]$.
Our framework incorporates noise into the diffusion process of each step of PPR. The privacy guarantees for this noisy PPR are derived from Theorem 4 with $\rho_{\text{diff}}=2\beta\eta$ and $\gamma_{\text{max}}=\beta$. Note that, since $\gamma_{j,k}>0$ for all $j\in\{1,2,3\}$ in PPR scenarios, all signals propagating among nodes should be non-negative. Consequently, the degree-based thresholding function $f$ can be further modified as $f(\mathbf{x})=\min(\max(\mathbf{x},\mathbf{0}),\eta\cdot\tilde{\mathbf{d}})$.